OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 4)(2x + 1)(x – 7)?

jimthompson5910 (jim_thompson5910):

f(x) = (x + 4)(2x + 1)(x – 7) 0 = (x + 4)(2x + 1)(x – 7) (x + 4)(2x + 1)(x – 7) = 0 x+4 = 0, 2x+1=0, or x-7=0 Solve each equation for x to get your three answers

OpenStudy (anonymous):

Thank you!!!

jimthompson5910 (jim_thompson5910):

np